# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v1

@@ Abstract
Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL). Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value. Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces. To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs). Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation. A custom-designed coder-evaluator framework is further employed to guide the generation, verification, and refinement of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference. Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios.
123
1 Introduction
@@
Recent advances in cooperative multi-agent reinforcement learning (MARL) [23, 3] have led to widespread applications in domains such as assembly lines [13], logistics distribution [42], autonomous driving [49], and swarm robotics [24]. In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training. A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called "lazy agents" [18].
@@
To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community. Attention-based credit assignment methods [33, 22] incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions. By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions. This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
@@
Policy gradient approaches [1] tackle credit assignment from a differentiable optimization perspective. Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability. These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation. In addition, value decomposition methods [35, 38] have become increasingly popular. These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent. The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution. These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
@@
However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for individual agents. A representative example of this is the mixing network employed in value decomposition methods. These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability [17, 29]. With the advancement of large language models (LLMs) [12], new paradigms have emerged to address many long-standing challenges in MARL. By leveraging extensive pre-trained knowledge [25], superior generalization capacity[2], and powerful code generation abilities [39], LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines. Recent studies have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals. Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines. Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency [30]. In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality [48]. Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
@@
In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network. To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation. Within this framework, LLMs act as both a code generator and a code evaluator [5]. Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT [28, 34, 10]. Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability. Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues [21] commonly encountered by LLMs. This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network. Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks.
The main contributions of this study are summarized as follows:
@@
We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF. This framework is generalizable and can be extended to other reinforcement learning scenarios enhanced by LLMs.
@@
2 Related Work
2.1 Credit Assignment in MARL
@@
In recent years, several algorithms have been proposed to address the credit assignment problem. And policy-based approaches [45, 47] were among the earliest introduced. MADDPG [19] extends the DDPG [15] framework to multi-agent settings by incorporating centralized training with decentralized execution. This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments. MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA [6] leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG. MAAC [11] further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
@@
Another widely adopted paradigm is the value decomposition approach [28, 40, 27, 41], which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN [35] models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX [28] improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN [32] and QPLEX [37] mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. LICA [50] adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information. Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks. RIIT [10] goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training. Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
2.2 LLM-enhanced RL
@@
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning [46, 14]. In the SMAC-R1 framework [4], agents utilize LLMs to generate decision tree code by providing task descriptions. The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards. Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments. They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes. For instance, the Reflexion framework [31] introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters. Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions. Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks. The EUREKA framework [20] achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
@@
These examples demonstrate the powerful information processing and code generation capabilities of LLMs. Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL. Qu et al. [26] were the first to introduce LLMs into the credit assignment problem. They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly. This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms. However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components. Additionally, Lin et al. [16] employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions. These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance. Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
3 Preliminaries
3.1 Decentralized Partially Observable Markov Decision Process
@@
A key challenge in Dec-POMDPs lies in the credit assignment problem. Since all agents share the same reward, it is difficult to attribute individual contributions. Moreover, as in other multi-agent settings, the environment is non-stationary from each agent’s perspective due to the concurrent learning processes of other agents, posing significant challenges to coordination and decentralized control.
3.2 Value Decomposition Methods
@@
Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions of individual agents into a global Q-value function . The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint. The IGM principle [9] requires that the joint action that maximizes the global action-value function must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
@@
where represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this, Qatten [43] provides a theoretical extension, showing that can be expressed as a linear combination of the individual functions:
@@
denotes a coefficient that depends on the state , where is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of with respect to , where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on is sufficiently small to be ignored or effectively absorbed into a bias term , allowing for a simplified model formulation and reduced computational complexity. Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
4 Method
@@
This section presents a detailed discussion of the implementation of each component of QLLM. We first introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination [36] and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
4.1 Training-Free Credit Assignment Function
@@
As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions. The network takes as input the individual Q-values of all agents, denoted as , and produces the joint Q-value as output. To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism [8]. Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state . This design facilitates state-aware adaptation while ensuring compatibility with the centralized training and decentralized execution (CTDE) framework, making deployment straightforward. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
@@
To address these issues, we propose Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM. This module is generated by LLMs through iterative error correction, based on a series of task-related prompts. As illustrated in Equation (2), the global Q-value function can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF. Its mathematical expression is as follows:
The coefficient is a state-dependent weight associated with the global state , and is a bias term that also depends on the global state. Both terms constitute key components of the TFCAF, which are generated by LLMs. In the following, we provide a detailed explanation of how the TFCAF is generated and refined through error correction, as well as how it facilitates credit assignment in multi-agent systems.
4.2 Coder-Evaluator Framework
Due to the susceptibility of LLMs to hallucinations, the generated TFCAF must undergo continuous refinement. To address this challenge, we propose the coder-evaluator framework, which automates the generation of TFCAF through a modular and iterative process, ensuring both accuracy and robustness. This framework employs two LLMs with distinct yet complementary roles: the coder LLM , responsible for generating candidate TFCAFs, and the evaluator LLM , which critically evaluates these candidates and provides constructive feedback for improvement. The interaction between the two models is driven by two types of prompts: task prompts, which describe the reinforcement learning task in detail, and role prompts, which define the specific responsibilities and operational roles of each LLM within the framework. The role prompts are further categorized into coder prompts and evaluator prompts, corresponding to the Coder LLM and the Evaluator LLM, respectively. The iterative generation and refinement of TFCAF can be formally described through the following three clearly-defined steps.

### v4

@@ Abstract
Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL). Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value. Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces. To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs). Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation. A custom-designed coder-evaluator framework is further employed to guide the generation and verification of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference. Furthermore, an IGM-Gating Mechanism enables QLLM to flexibly enforce or relax the monotonicity constraint depending on task demands, covering both IGM-compliant and non-monotonic scenarios. Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios. The code is available at https://github.com/zhouyangjiang71-sys/QLLM.
Introduction
@@
Recent advances in cooperative multi-agent reinforcement learning (MARL)(Oroojlooy and Hajinezhad 2023) have led to widespread applications in domains such as assembly lines (Kaven et al. 2024), logistics distribution (Yan et al. 2022), autonomous driving (Zhang et al. 2024b), and swarm robotics (Orr and Dutta 2023). In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training. A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called ”lazy agents” (Liu et al. 2023).
@@
To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community. Attention-based credit assignment methods (Soydaner 2022) incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions. By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions. This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
@@
Policy gradient approaches (Agarwal et al. 2021) tackle credit assignment from a differentiable optimization perspective. Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability. These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation. In addition, value decomposition methods (Sunehag et al. 2017) have become increasingly popular. These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent. The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution. These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
@@
However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for each agents. A representative example of this is the mixing network employed in value decomposition methods. These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability (Linardatos, Papastefanopoulos, and Kotsiantis 2020). With the advancement of large language models (LLMs) (Kasneci et al. 2023), new paradigms have emerged to address many long-standing challenges in MARL. By leveraging extensive pre-trained knowledge (Petroni et al. 2019), superior generalization capacity (Anil et al. 2022), and powerful code generation abilities (Xu et al. 2022a), LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines. Recent studies (Qu et al. 2025) have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals. Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines. Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency (Sanh et al. 2021). In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality (Zhang et al. 2023b). Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
@@
In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network. To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation. Within this framework, LLMs act as both a code generator and a code evaluator (Desmond et al. 2024). Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT (Rashid et al. 2020b; Su, Adams, and Beling 2021; Hu et al. 2021). Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability. Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues (Martino, Iannelli, and Truong 2023) commonly encountered by LLMs. This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network. Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks. An IGM-Gating Mechanism is further introduced to flexibly switch the monotonicity constraint on or off, adapting to tasks where the Individual-Global-Max principle holds or does not. The main contributions of this study are summarized as follows:
@@
We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF. This framework is generalizable and can be extended to other MARL scenarios enhanced by LLMs.
@@
Related Work
Credit Assignment in MARL
@@
In recent years, the credit assignment problem has emerged as a fundamental challenge in multi-agent systems. And policy-based approaches (Yu et al. 2022; Zhang et al. 2024a) were among the first to introduce credit assignment mechanisms. MADDPG (Lowe et al. 2017) extends the DDPG (Lillicrap et al. 2015) framework to multi-agent settings by incorporating centralized training with decentralized execution. This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments. MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA (Foerster et al. 2018) leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG. MAAC (Iqbal and Sha 2019) further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
@@
Another widely adopted paradigm is the value decomposition approach (Rashid et al. 2020b; Xu et al. 2022b), which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN (Sunehag et al. 2017) models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX (Rashid et al. 2020b) improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN (Son et al. 2019) and QPLEX (Wang et al. 2020) mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. LICA (Zhou et al. 2020) adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information. Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks. RIIT (Hu et al. 2021) goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training. Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
LLM-enhanced Reinforcement Learning
@@
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning (Zhang et al. 2023a; Li et al. 2024). In the SMAC-R1 framework (Deng et al. 2024), agents utilize LLMs to generate decision tree code by providing task descriptions. The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards. Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments. They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes. For instance, the Reflexion framework (Shinn et al. 2023) introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters. Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions. Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks. The EUREKA framework (Ma et al. 2023) achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
@@
These examples demonstrate the powerful information processing and code generation capabilities of LLMs. Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL. Qu et al. (2025) were the first to introduce LLMs into the credit assignment problem. They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly. This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms. However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components. Additionally, Lin et al. (2025) employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions. These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance. Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
Preliminaries
Decentralized Partially Observable Markov Decision Process
@@
A key challenge in Dec-POMDPs lies in the credit assignment problem. This issue arises because all agents within the system share a common reward structure, making it inherently difficult to accurately attribute individual contributions to the collective outcome. Furthermore, the environment is perceived as non-stationary from the viewpoint of each agent. This non-stationarity poses significant challenges in achieving effective coordination and optimal decision-making, as agents must adapt to the changing strategies of others while managing incomplete state representations.
Value Decomposition and IGM Principle
@@
Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions of individual agents into a global Q-value function . The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint. The IGM principle (Hong, Jin, and Tang 2022) requires that the joint action that maximizes the global action-value function must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
@@
where represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this, Qatten (Yang et al. 2020) provides a theoretical extension, showing that can be expressed as a linear combination of the individual functions:
@@
Here, denotes a coefficient that depends on the state , is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of with respect to , where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on is sufficiently small to be ignored or effectively absorbed into a bias term , allowing for a simplified model formulation and reduced computational complexity. Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
Method
@@
This section presents a detailed discussion of the implementation of each component of QLLM. We first propose the concept of the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and more relaxed joint value modeling. This provides a foundation for adaptively handling different types of multi-agent tasks. Then we introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination (Tonmoy et al. 2024) and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
IGM-Gating Mechanism
To flexibly handle diverse coordination demands across tasks, we introduce the IGM-Gating Mechanism, which enables the learning architecture to switch between strict adherence to the IGM principle and relaxed joint value modeling. This mechanism allows for conditional enforcement of the monotonicity constraint in the credit assignment.
Definition 1 (IGM-Gating Mechanism).
Let be the joint action-value function, and the individual utility functions. Define a binary gating variable , where indicates that IGM is enforced. The IGM-Gating Mechanism imposes the following conditional monotonicity constraint:
When the gating variable is inactive (), no monotonicity constraint is enforced, allowing to freely capture non-monotonic dependencies.
Training-Free Credit Assignment Function
@@
As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions. The network takes as input the individual Q-values of all agents, denoted as , and produces the joint Q-value as output. To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism (Ha, Dai, and Le 2017). Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state . This design facilitates state-aware adaptation while ensuring compatibility with CTDE framework, making deployment straightforward. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
@@
To address these issues, we propose the Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM. This module is generated by LLMs through iterative error correction, based on a series of easily accessible task-related prompts. As illustrated in Equation (2), the global Q-value function can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF, which we formally define below.
Definition 2.
A Training-Free Credit Assignment Function (TFCAF) is defined as:
where and are general nonlinear functions of the global state generated by LLMs before training.
As illustrated in Definition 1, when the gating signal , the function is strictly constrained to be non-negative. In contrast, when , this constraint is lifted. Benefiting from the rich prior knowledge of LLMs, the appropriate value of can be easily inferred from the task description, enabling the model to adapt to different types of tasks. For instance, in matrix game scenarios, where individual optimal decisions do not necessarily align with the global optimum, the LLM may assign a negative value to . The coefficients and the bias term make the overall TFCAF a nonlinear mapping, although its final combination with the local remains linear in form. Both components are automatically produced by LLMs through the coder-evaluator framework, which we detail next along with the iterative correction and its role in credit assignment for multi-agent systems.
Coder Evaluator Framework
Due to the susceptibility of LLMs to hallucinations, the process of generating TFCAF must be designed to be robust. To address this challenge, we propose the coder-evaluator framework, which automatically generates TFCAF through a modular and self-correcting process, ensuring both accurate credit assignment and logical consistency in code generation.
This framework employs two LLMs with distinct yet complementary roles: the coder LLM , responsible for generating candidate TFCAFs, and the evaluator LLM , which critically evaluates these candidates. All of these are generated by LLMs through structured prompts without any need for human intervention.
The interaction between the two models is driven by two types of prompts: task prompts () specify the meanings of each dimension in the global state space, the action space of the agents, and the composition of the reward function. This information is objective and can be easily obtained from the environment; role prompts () define the fixed responsibilities and operational roles of each LLM within the framework. The role prompts are further categorized into coder prompts and evaluator prompts, which remain unchanged across different RL tasks and correspond to the Coder LLM and the Evaluator LLM, respectively. This clear separation and stability of prompt design make QLLM easy to adapt to new cooperative tasks without extensive manual prompt tuning or subjective rewriting (see in Appendix A).
The generation of TFCAF can be formally described through the following three clearly-defined steps.